OpenStudy (anonymous):

Find the slope of the line passing through (4, 9) and (-2, 9). 0 undefined

OpenStudy (dtan5457):

slope formula if you remember:\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

yeh i remember

OpenStudy (dtan5457):

If you plot those points, does it make a vertical or horizontal line?

OpenStudy (dtan5457):

if it makes a vertical line: undefined horizontal:slope of 0

OpenStudy (anonymous):

i got -6for the slope

OpenStudy (dtan5457):

Nope. The slope is definitely either 0 or undefined. y2-y1=9-9=o/anything=0 or undefined

OpenStudy (dtan5457):

if it makes a vertical line: undefined horizontal:slope of 0 or if the x values change, it will be horizontal.

OpenStudy (dtan5457):

does it change here?

OpenStudy (anonymous):

0

OpenStudy (dtan5457):

it is indeed a horizontal line, which makes the slope 0. good job.

OpenStudy (anonymous):

thank you